Momney:

Which equation is equivalent to –k + 0.03 + 1.01k = –2.45 – 1.81k

Momney:

nsb

shisui:

man shadows been typing for a while

Shadow:

\[–k + 0.03 + 1.01k = –2.45 – 1.81k\] I added -k with 1.01k, think of k as 1.00k \[0.03 + 0.01k = -2.45 - 1.81k\] add 1.81 to both sides \[0.03 + 1.82k = -2.45\] Subtract 0.03 from both sides \[1.82k = -2.48\] Divide 1.82k form both sides \[k = \frac{ -2.48 }{ 1.82 } \rightarrow k = -1.36\] I just solved for k. Odds are one of equations will follow one of the above forms.

shisui:

i see why...

Shadow:

lol